List of Slides

1

– General Linux 1 –

Setup different PC expansion cards [2]

(Linux Professional Institute Certification)

a

   .~.
   /V\      by: geoffrey robertson
  // \\         geoffrey@zip.com.au
  @._.@
$Id: gl1.101.5.slides.tex,v 1.2 2003/05/30 04:58:40 waratah Exp $

aCopyright \begin{matrix}c \\ ◯ \end{matrix} 2002 Geoffrey Robertson. Permission is granted to make and distribute verbatim copies or modified versions of this document provided that this copyright notice and this permission notice are preserved on all copies under the terms of the GNU General Public License as published by the Free Software Foundation—either version 2 of the License or (at your option) any later version.

Setup different PC expansion cards [2]

(1.1) 101 Hardware & Architecture [11]

2.101.1 Configure BIOS Fundamental Settings [3] 2.101.2 Setup SCSI devices [2] 2.101.3

Setup different PC expansion cards [2] 2.101.4 Configure communication devices [2] 2.101.5 Set Up USB devices [2]

Setup different PC expansion cards [2]

Objective

Setup different PC expansion cards [2]

Key files, terms, and utilities

Setup different PC expansion cards [2]

Resources of interest The End

7